OpenStudy (blue_tiger):

use the information given for the year 1995 and 2000, letting x=5 represent 1995, x=10 represent 2000, and y the number of post offices, to write an equation that models the data. Graph (amount of post offices)- 1995 has 28,392, 2000 has 27,876

OpenStudy (moldybubblegum11):

Im sorry i dont know the answer to this but @563blackghost and @mathmate might be able to figure it out. D: sorry

OpenStudy (blue_tiger):

alright thanks anyway

OpenStudy (blue_tiger):

@Zarkon

OpenStudy (jerry45):

@zepdrix

OpenStudy (jerry45):

@Directrix

OpenStudy (jerry45):

@welshfella

OpenStudy (jerry45):

@mathmate

OpenStudy (jerry45):

wow no one came try PA bro

OpenStudy (blue_tiger):

PA?

OpenStudy (jerry45):

peer-answer basically the rivals of openstudy

OpenStudy (seratul):

I can help. Did you make a graph yet?

OpenStudy (seratul):

Just the outline of it.

OpenStudy (blue_tiger):

no, it hasn't said that i have to graph it

OpenStudy (seratul):

Oh okay, one second.

OpenStudy (seratul):

Sorry, I have no clue what the question is asking. The wording is confusing.

OpenStudy (mathmate):

There are many ways to model a set of data, linear, exponential, square, sqaure-root. With only two points \(any\)thing goes! If it is linear, then use y=ax+b If it is exponential, try y=ab^x If it is square, try y=ax^2+bx+c (but you have to assume one of a, b or c) and so on. Find the unknown constants using the given data and you'll have it!